OpenStudy (anonymous):

write in radian form using pi: 125 degrees

zepdrix (zepdrix):

There are 2pi radians in one rotation around the unit circle, or 360 degrees. In half a circle we can think just take half of those distances. (pi or 180 degrees). We generally use this for converting from radians to degrees, or the other way around. \[\huge 125^o\cdot \left(\frac{\pi}{180^o}\right)\] We put degrees on the bottom, so they'll cancel out with the degrees on the top.

zepdrix (zepdrix):

Then its just a matter of simplifying things down :D \[\huge \frac{125 \pi}{180}\]

zepdrix (zepdrix):

Confused about any of that? :D

OpenStudy (anonymous):

i know its 2.18 but i dont know how to write that value using pi

zepdrix (zepdrix):

That's what I'm trying to explain :3 You want to multiply your quantity by THIS fraction when converting from degrees to radians\[\huge \left(\frac{\pi}{180}\right)\] And use THIS fraction when converting from radians to degrees \[\huge \left(\frac{180}{\pi}\right)\]

zepdrix (zepdrix):

\[\huge 125\cdot \left(\frac{\pi}{180}\right)=\left(\frac{125\pi}{180}\right)\]Simplifying it down can be a little tricky, you just need to look for numbers that will work. Hmmm it looks like we can factor a 5 out of both of them.

zepdrix (zepdrix):

\[\huge \left(\frac{25\cdot 5 \pi}{36\cdot 5}\right)=\left(\frac{25\pi}{36}\right)\]

OpenStudy (anonymous):

ohh i see......

OpenStudy (anonymous):

thanks!